OpenStudy (anonymous):

Find the average value of the function f on the given interval. f(x)=1/x^2 on the interval [1,5] I'll draw what I've got so far.

OpenStudy (abb0t):

Take the integral of x^-2 from 1<x<5

OpenStudy (anonymous):



OpenStudy (anonymous):

What is the next step?

OpenStudy (abb0t):

\[\frac{ 1 }{ 4 }\left[ -\frac{ 1 }{ 5 } - \left( -\frac{ 1 }{ 1 } \right) \right]=\]

OpenStudy (anonymous):

Thank you that helps me a bunch! So the answer is 1/5 and now I know the

OpenStudy (anonymous):

process.

OpenStudy (abb0t):

Read up on the fundamental theorem of Calculus II. It explains this clearly.